# Learning Tractable Distributions of Language Model Continuations

### v1

@@ Abstract
Controlled language generation conditions text on sequence-level constraints (for example, syntax, style, or safety). These constraints may depend on future tokens, which makes directly conditioning an autoregressive language model (LM) generally intractable. Prior work uses tractable surrogates such as hidden Markov models (HMMs) to approximate the distribution over continuations and adjust the model’s next-token logits at decoding time. However, we find that these surrogates are often weakly context aware, which reduces query quality. We propose Learning to Look Ahead (LTLA), a hybrid approach that pairs the same base language model for rich prefix encoding with a fixed tractable surrogate model that computes exact continuation probabilities. Two efficiency pitfalls arise when adding neural context: (i) naïvely rescoring the prefix with every candidate next token requires a sweep over the entire vocabulary at each step, and (ii) predicting fresh surrogate parameters for each prefix, although tractable at a single step, forces recomputation of future probabilities for every new prefix and eliminates reuse. LTLA avoids both by using a single batched HMM update to account for all next-token candidates at once, and by conditioning only the surrogate’s latent state prior on the LM’s hidden representations while keeping the surrogate decoder fixed, so computations can be reused across prefixes. Empirically, LTLA attains higher conditional likelihood than an unconditional HMM, approximates continuation distributions for vision–language models where a standalone HMM cannot encode visual context, and improves constraint satisfaction at comparable fluency on controlled-generation tasks, with minimal inference overhead.
@@ 1 Introduction
Autoregressive models are the dominant way to represent high-dimensional discrete distributions over language and sequence data (grattafiori2024llama; shin2021protein), factorizing the distribution into a sequence of next-token conditional distributions. Many useful queries, however, concern properties of the entire sequence: to generate autoregressively under a sequence-level constraint, we need the probability that the remaining suffix will satisfy that constraint given the current prefix (boyd2022predictive; ZhangICML23). Computing this probability requires reasoning over exponentially many possible futures and is therefore intractable for standard autoregressive language models (LM). In practice, estimates come from sampling (qin2022cold; hu2023amortizing; lew2023sequential), which introduces significant computational overhead, or from learned heuristics (krause2020gedi; yang2021fudge; meng2022controllable) that must be tailored to each query or constraint.
Tractable probabilistic models (TPMs) are a broad class of generative models that admit efficient routines for exactly computing such conditional queries exactly and efficiently (ProbCirc20). As such, a natural approach employed by prior works (ZhangICML23; YidouWengICML25) is to utilize TPMs as tractable surrogates to the LM, such that queries on the LM can be approximated by computing them on the surrogate. However, we identify that current TPMs used for language modeling, such as hidden Markov models (HMM), face two significant obstacles in this setting. Firstly, we find empirically that HMMs are comparatively insensitive to the prefix, which yields less accurate distributions over continuations, as illustrated in Fig. 1. Secondly, as models operating only over discrete sequences, they cannot effectively encode some kinds of context, notably continuous multimodal features used by modern vision–language models.
To encode the past well and answer future queries quickly, we separate the two jobs. Understanding the past (“lookback”) should preserve as much information as possible, whereas forecasting the future (“lookahead”) must remain tractable. The same base LM already builds a rich representation of the prefix, which we reuse to encode the prefix into the tractable model (for example, a HMM) that models and computes exact probabilities over continuations. In short, we propose a hybrid tractable model where a transformer-based LM is used for lookback and a HMM for lookahead.
@@
Figure 1: The encoder given by standard HMMs is often insensitive to information contained within the context. In this example, we show an example with the context they fired the <x> after just one, where <x> can be coach or employee. The distribution is almost identical for the standard HMM, while the neural HMM shows a significant shift in distribution (in particular, with season and game being more likely when <x> = coach).
@@
Making this hybrid model practical raises two concrete challenges. First, when autoregressively generating the next token, a naïve strategy would run the transformer on the prefix concatenated with every candidate token across the vocabulary to assess how each affects the downstream conditional query; this per-step sweep over the entire vocabulary is expensive (see “Exhaustive LM Rescoring” in Table 1). Second, conditioning the surrogate itself on the prefix by predicting fresh parameters for every context remains tractable at a single step but prevents reuse: as the prefix grows, the surrogate must be rebuilt and its future probabilities recomputed at every step, which increases decode-time cost and memory (see “Prefix-Parameterized TPM” in Table 1; cf. Sec. 3.1). Both issues undercut the efficiency that makes TPMs attractive.
@@
Thus, we propose a solution that encodes rich context with the transformer while maintaining TPM efficiency by encoding the prefix only into the HMM surrogate’s latent prior and keeping the rest of the HMM fixed. Specifically, the transformer’s hidden states are fed into to a lightweight head that is trained to produce a prior over the surrogate’s latent state, and the observed next token $x_{t}$ is incorporated with a one-step HMM update as a single batched matrix–vector operation, avoiding separate computation for each token in the vocabulary. With fixed transition and emission parameters, future computations are reusable across steps, so conditional queries remain exact and fast even for long sequence lengths (see “LTLA (Ours)” in Table 1). We call this approach Learning to Look Ahead (LTLA): it learns to injects context through the LM representation, while preserving tractability of the lookahead model.
@@
1. We introduce LTLA, a novel approach for learning tractable surrogates to language models. LTLA reuses the base language model to predict the prior of a HMM modeling continuations given the current context, enabling accurate estimates of conditional queries at each time step of generation while only using a small number of matrix-vector multiplications. We further show how LTLA can be used to condition on multimodal context by leveraging vision-language models (VLM).
2. We propose different architectural choices for both the neural encoder and tractable decoder, and investigate their effect on modeling performance, decoding overhead, and downstream performance.
3. Empirically, we show that LTLA achieves improved conditional log-likelihood compared with existing tractable models, especially for the next few tokens. We also demonstrate the application of LTLA to controlled-generation tasks for both language and vision–language models, where LTLA improves constraint satisfaction compared to existing approaches while adding only a small inference-time overhead.
@@ 2 Tractable Modeling of Sequences
In this work, we are interested in autoregressive sequence models, and how to effectively answer queries about the distribution over sequences that they represent. An autoregressive model decomposes the distribution over a sequence of tokens $x_{1:T}$ as follows:
$p(x_{1:T})=\prod_{t=1}^{T}p(x_{t}|x_{<t}).$ (1)
@@
Aside from generating or analyzing the next token distribution, we are often interested in more complex properties of the distribution. These can be represented generally as conditional probability queries (boyd2022predictive), which ask for the probability $p(\alpha|x_{1:t})$ of some event $\alpha$, where $x_{1:t}$ is the prefix (or context) generated so far. Examples of events $\alpha$ might include the $k^{\text{th}}$ token $x_{t+k}$ in the future taking a particular value, some token $a$ appearing before token $b$ in the sequence, the expected length of the sequence generated, or more complex properties involving grammatical or semantic constraints (ZhangNeurIPS24; YidouWengICML25; ahmed2025semantic). What these queries have in common is that they require looking into the future: that is, aggregating over all possible continuations, weighted by their conditional probability given the context:
@@
Consider the problem of estimating the conditional probability query $p(\alpha|x_{1:t})$ given in Equation 2. The key tradeoff is between (i) the accuracy and (ii) the computation cost of the estimation. Explicitly enumerating all such continuations would result in an exact answer, but is clearly infeasible as the number of such continuations grows exponentially with sequence length. As such, for most models, one typically needs to approximate, for example by (i) using sampling-based techniques targeting the conditional continuation distribution (qin2022cold; zhao2024probabilistic; loula2025syntactic); or (ii) directly approximating the conditional probability query using a neural classifier or generative model specialized to the constraint $\alpha$ (krause2021gedi; yang2021fudge; meng2022controllable).
In this work, we consider an alternative, computationally efficient approach based upon tractable modeling of continuations. This stems from the observation that, for certain distributions $p$ and queries $\alpha$, the computation of $p(\alpha|x_{1:t})$ can be done both (i) exactly and (ii) efficiently, breaking the tradeoff. In particular, tractable probabilistic models (TPMs) (ProbCirc20) are classes of probabilistic generative models that are known to enable computationally efficient analytical computation of many classes of queries, such as marginal probabilities.
Example 1
@@
$q(x_{1},\ldots,x_{T})=\sum_{z_{1},\ldots,z_{T}}q(z_{1})q(x_{1}|z_{1})\prod_{t=2}^{T}q(z_{t}|z_{t-1})q(x_{t}|z_{t}).$
HMMs enable efficient computation of various queries via (variants of) the forward and backward algorithms. For example, if $\alpha$ is the event that the last token $X_{T}=\texttt{world}$, we can compute the conditional query $q(\alpha|x_{1:t})$ using the fact that (by conditional independence):
@@
with base case $q(\alpha|z_{T})=q(x_{T}=\texttt{world}|z_{T})$. Prior work has also shown that HMMs support tractable querying of many other conditions $\alpha$, including complex logical constraints111more specifically, deterministic finite automata (DFA) or unambiguous context-free grammars (uCFG). (ZhangNeurIPS24; ZhangAISTATS25) or factorized classifiers for semantic constraints (YidouWengICML25).
@@
The distribution $p$ we are interested in will typically not be tractable in this way. However, we can aim to approximate the (prior) distribution $p(x_{t+1:T}|x_{1:t})$ using a simpler tractable surrogate model $q(x_{t+1:T}|x_{1:t})$ for which conditional queries are tractable, and estimating $p(\alpha|x_{1:t})\approx q(\alpha|x_{1:t})$ using the tractable approximation $q$. This approach has been employed in prior works (ZhangICML23) using HMMs as the TPM of choice, which has been shown to lead to state-of-the-art performance on controlled generation benchmarks.
The Promise and Challenges of Tractable Surrogate Modeling
@@
Besides state-of-the-art empirical performance on downstream applications, the tractable modeling approach offers a number of other benefits related to computational efficiency. Firstly, it amortizes the cost across different constraints $\alpha$. That is, since $q$ is trained to match the prior, and the computation of the query $q(\alpha|x_{1:t})$ is conducted using a symbolic algorithm, one does not need to commit in advance to any particular condition $\alpha$. This is in contrast to approaches that train models to specifically target the posterior $p(x_{t+1:n}|x_{1:t},\alpha)$. Secondly, tractable models can amortize the cost across different contexts $x_{1:t}$ by exploiting conditional independence. For example, in the HMM query computation in Equation 4, the backward quantity $q(\alpha|z_{t})$ can be precomputed and cached independently of the context $x_{1:t}$.
@@
This computational efficiency, however, also comes at a cost. The quality of the query estimate provided by the tractable model depends significantly on the quality of the approximation of $q$ to $p$, which in turn has been shown to affect downstream performance (ZhangICML23; YidouWengICML25). Unfortunately, the class of tractable models (e.g., HMMs) is fundamentally less expressive than e.g., neural autoregressive models (choi2019relative; BroadrickICML25). This is reflected in our observation in Figure 1, where a HMM trained to approximate a GPT2-large language model is unable to effectively encode dependence on context $x_{1:t}$ in its distribution $q(x_{t+1:n}|x_{1:t})$. As such, a key challenge is to improve the expressivity and learning performance of the tractable model approximation, while maintaining the computational efficiency of the existing HMM-based approximation. In the next section, we will present our approach, Learning To Look Ahead (LTLA), which utilizes an amortized inference approach with a neural encoder to obtain the tractable approximation $q$.
@@ 3 Learning to Look Ahead
Unfortunately, in practice we cannot afford to optimize $q^{(j)}$ separately for every context $x_{1:t}^{(j)}$; instead we must take an amortized inference approach, in which we learn to predict $q(x_{t+1:T})$ given context $x_{1:t}$. In particular, an HMM trained on the joint distribution over contexts and continuations can be viewed as performing amortized inference by (i) applying a probabilistic encoder $q(z_{t}|x_{1:t})\in\mathbb{R}^{H}_{\geq 0}$ to predict the latent state distribution and (ii) parameterizing the distribution over continuations via this latent state and the probabilistic decoder $q(x_{t+1:T}|z_{t})$. Our key insight is that, for answering queries about the continuation, we only need to be able to (i) evaluate the encoder and (ii) answer queries about the decoder’s distribution. As such, only the decoder needs to be tractable, and we can increase the expressivity of the continuation model by allowing the encoder to be an arbitrary neural network $q_{enc}$, giving rise to the following hybrid model:
@@
We call our method Learning To Look Ahead (LTLA) as (i) we train a neural encoder to learn a language model’s distribution over continuations, and (ii) we execute a symbolic algorithm to look ahead over exponentially many continuations with a probability distribution given by the output of the learned encoder. Although we only maintain the model’s tractability over the continuation tokens $x_{t+1:T}$, this is sufficient for downstream applications where tokens are generated autoregressively. In the rest of this section, we will first describe the algorithms for and complexity of conditional probability queries of interest, before discussing architectural choices for the neural encoder.
3.1 Inference using Hybrid HMMs
After conditioning on a context $x_{\leq t}$, we obtain a tractable model of the distribution of continuations $q(x_{>t}|x_{\leq t})$, rendering many natural queries exactly and efficiently computable. For example, after observing a sequence $x_{\leq t}$, we might ask the following questions: (i) what is the probability that the sequence ends within the next $k$ tokens; (ii) what is the probability that the sequence contains a particular keyword; (iii) what is the probability that the generated sequence will be labeled “toxic”/“biased”/etc by a given tractable classifier? In general, queries $q(\alpha|x_{\leq t})$, where $\alpha$ is some subset of the possible continuations $x_{>t}$, can be written $q(\alpha|x_{\leq t})=\sum_{x_{>t}}q(x_{>t}|x_{\leq t})\mathbbm{1}\{x_{>t}\in\alpha\}$ where $\mathbbm{1}\{x_{>t}\in\alpha\}$ is an indicator function of the membership of $x_{>t}$ in $\alpha$. So, to efficiently compute $q(\alpha|x_{\leq t})$, it suffices to be able to multiply $q(x_{>t}|x_{\leq t})$ by the indicator function for $\alpha$ in such a way that the summation remains tractable; general sufficient conditions are known for such tractable multiplications (darwiche2002knowledge; ZhangAISTATS25). More generally, tractable models also support other types of queries, including for example information-theoretic quantities like entropies and divergences (vergari2021compositional; wang2024compositional).

### v2

@@ Abstract
Controlled generation imposes sequence-level constraints (syntax, style, safety) that depend on future tokens, making exact conditioning of an autoregressive LM intractable. Tractable surrogates such as HMMs can approximate continuation distributions and steer decoding, but standard surrogates are often weakly context-aware. We propose Learning to Look Ahead (LTLA), a hybrid method that uses base-LM embeddings to condition a globally learned tractable surrogate: a neural head predicts only a prefix-dependent latent prior, while a shared HMM answers continuation queries exactly. LTLA is designed to avoid two common efficiency traps when adding neural context. First, it avoids vocabulary-sized prefix rescoring ($V$ extra LM evaluations) by scoring all next-token candidates via a single batched HMM forward update. Second, it avoids predicting a new HMM per prefix by learning one shared HMM and conditioning only the latent prior, which enables reuse of cached future-likelihood (backward) messages across decoding steps. Empirically, LTLA improves continuation likelihood over standard HMM surrogates, enables lookahead control for vision–language models by incorporating continuous context, achieves 100% syntactic constraint satisfaction, and improves detoxification while adding only a 14% decoding-time overhead.
keywords todo
@@ 1 Introduction
Autoregressive models are the dominant way to represent high-dimensional discrete distributions over language and sequence data (Grattafiori et al., 2024; Shin et al., 2021), factorizing the distribution into a sequence of next-token conditional distributions. Many useful queries, however, concern properties of the entire sequence: to generate autoregressively under a sequence-level constraint, we need the probability that the remaining suffix will satisfy that constraint given the current prefix (Boyd et al., 2022; Zhang et al., 2023). Computing this probability requires reasoning over exponentially many possible futures and is therefore intractable for standard autoregressive language models (LM). In practice, estimates come from sampling (Qin et al., 2022; Hu et al., 2023; Lew et al., 2023), which introduces significant computational overhead, or from learned heuristics (Krause et al., 2020; Yang and Klein, 2021; Meng et al., 2022) that must be tailored to each query or constraint.
@@
Figure 1: The encoder of standard HMMs is often insensitive to information contained within the context. For example, with the context they fired the <x> after just one, where <x> can be coach or employee, the distributions are almost identical for the standard HMM, while our neural HMM shows a significant shift in distribution (e.g, with season and game more likely for coach).
Tractable probabilistic models (TPMs) are a broad class of generative models that admit efficient routines for computing such conditional queries exactly (Choi et al., 2020). As such, a natural approach employed by prior works (Deng et al., 2025; Zhang et al., 2023, 2024; Yidou-Weng et al., 2025) is to utilize TPMs as tractable surrogates to the LM, such that queries on the LM can be approximated by computing them on the surrogate. However, TPMs commonly used for language modeling, such as hidden Markov models (HMMs), face two obstacles in this role. First, we find empirically that standard HMMs are often insensitive to prefix information, leading to less accurate continuation distributions. For example, in Fig. 1, changing <x> from coach to employee barely changes the HMM’s next-token distribution, despite the clear semantic change in the prefix. Second, as discrete sequence models, they do not naturally incorporate continuous context, such as image embeddings produced by modern vision–language models (VLMs).
Our key insight is that, to encode past context effectively and answer future queries efficiently, these two tasks should be decoupled. Understanding the past (lookback) should preserve as much information as possible, whereas forecasting the future (lookahead) must remain tractable. To achieve this, we reuse the base LM’s rich representation of the prefix to condition a tractable probabilistic inference model (e.g., an HMM) over continuations, yielding a lookahead distribution that is sensitive to semantic differences in the context (e.g., shifting mass toward season and game when <x>=coach in Fig. 1). In this work, we propose a hybrid tractable model that uses a transformer-based LM for lookback and an HMM for lookahead.
@@
Making this hybrid model practical raises two concrete challenges. First, when autoregressively generating the next token, a naïve strategy would run the transformer on the prefix concatenated with every candidate token across the vocabulary to assess how each affects the downstream conditional distribution; this per-step sweep over the entire vocabulary is expensive (see “Exhaustive LM Rescoring” in Table 1) and conceptually corresponds to a branching tree of continuations, as shown in Fig. 2(a). Second, conditioning the surrogate itself on the prefix by predicting fresh parameters for every context remains tractable at a single step but prevents reuse: as the prefix grows, the surrogate must be rebuilt and its future probabilities recomputed at every step, which increases decode-time cost and memory (see “Prefix-Parameterized TPM” in Table 1; cf. Sec. 3.1). Both issues undercut the efficiency that makes TPMs attractive.
@@
Thus, we encode rich context with the transformer while preserving TPM efficiency by conditioning only the HMM surrogate’s latent-state prior on the prefix, and using a single globally learned HMM shared across contexts. Specifically, the transformer’s hidden states are fed into a lightweight head that is trained to produce a prior over the surrogate’s latent state, and the observed next token $x_{t}$ is incorporated with a one-step HMM update as a single batched matrix–vector operation, avoiding separate computation for each token in the vocabulary. With fixed transition and emission parameters, future computations are reusable across steps, so conditional queries remain exact and fast even for long sequence lengths (see “LTLA” in Table 1). Fig. 2(b) illustrates this design: the LM handles lookback, while a linear-chain HMM surrogate performs lookahead with a single dynamic program. We call this approach Learning to Look Ahead (LTLA).
@@
1. We introduce LTLA, a hybrid approach for learning tractable surrogates to language models. LTLA reuses the base LM to predict a prefix-conditioned latent prior for an HMM over continuations, enabling accurate conditional queries at each generation step with only a small number of matrix–vector operations. This prefix conditioning makes the surrogate more context-sensitive, producing sharper shifts toward compatible continuations (e.g., Fig. 1), and it extends naturally to multimodal context in VLMs.
2. We study architectural choices for both the neural encoder and the tractable decoder, and quantify their trade-offs in surrogate accuracy and decoding overhead. In our experiments, a lightweight linear head already captures most of the gains while adding negligible overhead, so we use it as the default configuration.
3. Empirically, LTLA improves conditional log-likelihood over standard tractable surrogates, reducing perplexity by roughly 10–15% across hidden sizes. The largest gains occur in the next few tokens: next-token perplexity drops by 60%. On controlled generation, LTLA achieves 100% syntactic constraint satisfaction (e.g., CommonGen) and substantially improves semantic control in detoxification, reducing average toxicity from 0.254 to 0.015, while adding only a 14% inference-time overhead.
[…]
Figure 2: Naive LM lookahead vs LTLA. (a) Naive lookahead branches over many candidate continuations and requires a separate LM forward pass for each path, which is exponential as the horizon grows. (b) LTLA runs the LM once on the prefix to obtain a context embedding $h_{t}$, maps $h_{t}$ through a lightweight head to a prior over latent states $z_{t}$, and then uses an HMM to model future continuations. Lookahead scores are obtained by a single linear dynamic program instead of repeated LM calls.
@@ 2 Tractable Modeling of Sequences
We study autoregressive sequence models, and how to effectively answer queries about the distribution they represent. An autoregressive model decomposes the distribution over a sequence of tokens $x_{1:T}$ as $p(x_{1:T})=\prod_{t=1}^{T}p(x_{t}|x_{<t}).$
@@
Aside from generating or analyzing the next token distribution, we are often interested in more complex properties of the distribution. These can be represented generally as conditional probability queries (Boyd et al., 2022), which ask for the probability $p(\alpha|x_{1:t})$ of some event $\alpha$, where $x_{1:t}$ is the prefix (or context) generated so far. Examples of events $\alpha$ might include the $k^{\text{th}}$ token $x_{t+k}$ in the future taking a particular value, some token $a$ appearing before token $b$ in the sequence, the expected length of the sequence generated, or more complex properties involving grammatical or semantic constraints (Zhang et al., 2024; Yidou-Weng et al., 2025; Ahmed et al., 2025; Deng et al., 2025). What these queries have in common is that they require looking into the future: that is, aggregating over all possible continuations, weighted by their conditional probability given the context:
@@
Consider the problem of estimating the conditional probability query $p(\alpha|x_{1:t})$ given in Equation 1. The key tradeoff is between (i) the accuracy and (ii) the computation cost of the estimation. Explicitly enumerating all such continuations would result in an exact answer, but is clearly infeasible as the number of such continuations grows exponentially with sequence length. As such, for most models, one typically needs to approximate, for example by (i) using sampling-based techniques targeting the conditional continuation distribution (Qin et al., 2022; Zhao et al., 2024; Loula et al., 2025); or (ii) directly approximating the conditional probability query using a neural classifier or generative model specialized to the constraint $\alpha$ (Krause et al., 2021; Yang and Klein, 2021; Meng et al., 2022).
In this work, we consider an alternative, computationally efficient approach based upon tractable modeling of continuations. This stems from the observation that, for certain distributions $p$ and queries $\alpha$, the computation of $p(\alpha|x_{1:t})$ can be done both (i) exactly and (ii) efficiently, breaking the tradeoff. In particular, tractable probabilistic models (TPMs) (Choi et al., 2020) are classes of probabilistic generative models that are known to enable computationally efficient analytical computation of many classes of queries, such as marginal probabilities.
Example 1.
@@
$q(x_{1:T})=\sum_{z_{1},\ldots,z_{T}}q(z_{1})q(x_{1}|z_{1})\prod_{t=2}^{T}q(z_{t}|z_{t-1})q(x_{t}|z_{t}).$
HMMs enable efficient computation of various queries via (variants of) the forward and backward algorithms (Rabiner and Juang, 1986). For example, if $\alpha$ is the event that the last token $X_{T}=\texttt{world}$, we can compute the conditional query $q(\alpha|x_{1:t})$ using the fact that (by conditional independence):
@@
with base case $q(\alpha|z_{T})=q(x_{T}=\texttt{world}|z_{T})$. Prior work has also shown that HMMs support tractable querying of many other conditions $\alpha$, including complex logical constraints [^1] Specifically, constraints represented as deterministic finite automata (DFA) or unambiguous context-free grammars (uCFG). (Zhang et al., 2024, 2025b; Marzouk and de La Higuera, 2022; Deng et al., 2025) or factorized classifiers for semantic constraints (Vergari et al., 2021; Yidou-Weng et al., 2025).
@@
The LLM distribution $p$ we are interested in will not be tractable in this way. However, we can aim to approximate the (prior) distribution $p(x_{t+1:T}|x_{1:t})$ using a simpler tractable surrogate model $q(x_{t+1:T}|x_{1:t})$ for which conditional queries are tractable, and estimating $p(\alpha|x_{1:t})\approx q(\alpha|x_{1:t})$ using the tractable approximation $q$. This approach has been employed in prior works (Deng et al., 2025; Zhang et al., 2024; Yidou-Weng et al., 2025) using HMMs as the TPM of choice, and has been shown to lead to state-of-the-art performance on controlled generation benchmarks.
Promise and Challenges of Tractable Surrogates
@@
Besides state-of-the-art empirical performance on downstream applications, the tractable modeling approach offers a number of other benefits related to computational efficiency. Firstly, it amortizes the cost across different constraints $\alpha$. That is, since $q$ is trained to match the prior, and the computation of the query $q(\alpha|x_{1:t})$ is conducted using a symbolic algorithm, one does not need to commit in advance to any particular condition $\alpha$. This is in contrast to approaches that train models to specifically target the posterior $p(x_{t+1:n}|x_{1:t},\alpha)$. Secondly, tractable models can amortize the cost across different contexts $x_{1:t}$ by exploiting conditional independence. For example, in the HMM computation in Equation 3, the backward quantity $q(\alpha|z_{t})$ can be precomputed and cached independently of the context $x_{1:t}$.
@@
This computational efficiency, however, also comes at a cost. The quality of the query estimate provided by the tractable model depends significantly on the quality of the approximation of $q$ to $p$, which in turn has been shown to affect downstream performance (Zhang et al., 2023; Yidou-Weng et al., 2025; Deng et al., 2025). Unfortunately, the class of tractable models (e.g., HMMs) is fundamentally less expressive than e.g., neural autoregressive models (Choi et al., 2019; Broadrick et al., 2025). This is reflected in our observation in Figure 1, where a HMM trained to approximate a GPT2-large language model is unable to effectively encode dependence on context $x_{1:t}$ in its distribution $q(x_{t+1:n}|x_{1:t})$. As such, a key challenge is to improve the expressivity and learning performance of the tractable model approximation, while maintaining the computational efficiency of the existing HMM-based approximation. In the next section, we will present our approach, Learning To Look Ahead (LTLA), which utilizes an amortized inference approach with a neural encoder to obtain the tractable approximation $q$.
@@ 3 Learning to Look Ahead
Unfortunately, in practice we cannot afford to optimize $q^{(j)}$ separately for every context $x_{1:t}^{(j)}$; instead we must take an amortized inference approach, in which we learn to predict $q(x_{t+1:T})$ given context $x_{1:t}$. In particular, an HMM trained on the joint distribution over contexts and continuations can be viewed as performing amortized inference by (i) applying a probabilistic encoder $q(z_{t}|x_{1:t})$ to predict the latent state distribution and (ii) parameterizing the distribution over continuations via this latent state and the probabilistic decoder $q(x_{t+1:T}|z_{t})$. Our key insight is that, for answering queries about the continuation, we only need to be able to (i) evaluate the encoder and (ii) answer queries about the decoder’s distribution. As such, only the decoder needs to be tractable, and we can increase the expressivity of the continuation model by allowing the encoder to be an arbitrary neural network $q_{enc}$, giving rise to the following hybrid model:
@@
We call our method Learning To Look Ahead (LTLA) as (i) we train a neural encoder to learn a language model’s distribution over continuations, and (ii) we execute a symbolic algorithm to look ahead over exponentially many continuations with a probability distribution given by the output of the learned encoder. Next, we describe the algorithms for and complexity of conditional probability queries of interest, and then discuss architectural choices for the neural encoder.
Figure 3: Neural architectures for neural HMMs: (a) frozen Transformer with linear mapping, (b) frozen Transformer with additional learnable layer, (c) fully finetuned Transformer.
3.1 Inference using Neural HMMs
After conditioning on a context $x_{\leq t}$, we have a tractable model of the distribution of continuations $q(x_{>t}|x_{\leq t})$, rendering many natural queries exactly and efficiently computable. For example, we might ask (i) what is the probability that the sequence ends within the next $k$ tokens; (ii) what is the probability that the sequence contains a particular keyword; (iii) what is the probability that the generated sequence will be labeled “toxic”/“biased”/etc by a given tractable classifier? In general, queries $q(\alpha|x_{\leq t})$, where $\alpha$ is some subset of the possible continuations $x_{>t}$, can be written $q(\alpha|x_{\leq t})=\sum_{x_{>t}}q(x_{>t}|x_{\leq t})\mathbbm{1}\{x_{>t}\in\alpha\}$ where $\mathbbm{1}\{x_{>t}\in\alpha\}$ is an indicator function of the membership of $x_{>t}$ in $\alpha$. So to compute $q(\alpha|x_{\leq t})$, it suffices to multiply $q(x_{>t}|x_{\leq t})$ by the indicator function for $\alpha$ in such a way that the summation remains tractable; general sufficient conditions are known for such tractable products (Darwiche and Marquis, 2002; Zhang et al., 2025b). More generally, tractable models can support other types of queries, including for example information-theoretic quantities (Vergari et al., 2021; Wang et al., 2024).